[Hallc_running] RC Update

Dipangkar Dutta ddutta07 at gmail.com
Fri Jan 12 09:00:18 EST 2018

Hi All,
Yesterday the Hall was locked up soon after the DOE Under-secretary's tour
of the Hall. Last night over the swing and owl shifts MCC was able to
configure Halls A and B and send beam to all 3 Halls (D,A, B) at 5 pass.
This means that today during the day shift they will be setting up Hall C
for 3 pass beam. If all goes well we expect to see beam in the Hall during
the day shift.
Dave Meekins is already working on condensing hydrogen in the target.
SHMS is at 8 deg, ready to start the beam steering test as soon as beam is

-------------- next part --------------
An HTML attachment was scrubbed...
URL: <https://mailman.jlab.org/pipermail/hallc_running/attachments/20180112/dcaf9c2a/attachment.html>

More information about the Hallc_running mailing list